OpenStudy (anonymous):

what is the equation that makes this one (T=2pi sqrt(m/k)) be equal to k?

OpenStudy (mrdoe):

\[2\pi \sqrt{m/k} = m/k=(2\pi)^2 = m=(2\pi)^2*k = k=\frac{ m }{ (2\pi)^2 }\]

OpenStudy (mrdoe):

provided i read your equation correctly

OpenStudy (anonymous):

where did the time go? hold on let me use this equation button

OpenStudy (mrdoe):

ah sorry missed the T, just carry it through with the steps i showed you and you'll be fine

OpenStudy (anonymous):

\[T=2\pi \sqrt{m/k}\]

OpenStudy (anonymous):

I need this equation to solve for K so...k=

OpenStudy (mrdoe):

start by getting rid of the square root, T^2=(2pi)^2*(m/k) just some algebra from there

OpenStudy (mrdoe):

\[\frac{ T^2 }{ (2\pi)^2 }=m/k = ...\]

OpenStudy (anonymous):

what do you do to get m on the other side?

OpenStudy (mrdoe):

you will have to multiply both sides by k and then solve for k

OpenStudy (anonymous):

Thank you so much!

OpenStudy (mrdoe):

yvw